OpenStudy (anonymous):

The probability that a train leaves on time is 0.8. The probability that the train arrives on time and leaves on time is 0.24. What is the probability that the train arrives on time given that it leaves on time?

OpenStudy (anonymous):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}\]\

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

sooo .24/.8? because a&B would be B right?

OpenStudy (anonymous):

\[A\cap B\] is A and B, i.e arrive on time and leave on time, so \(P(A\cap B)=.24\) yes

OpenStudy (anonymous):

wait so it would be .24/.24? o.o;

OpenStudy (anonymous):

and \(P(B)=.8\) because you are told it is so yes, \[\frac{.24}{.8}=\frac{24}{80}=\frac{3}{10}\]

OpenStudy (anonymous):

oh no you were right the first time

OpenStudy (anonymous):

ohhh okay! Thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

quite a mixture of problems must be an on line class little of this little of that

OpenStudy (anonymous):

Yeah it's a review for Algebra 2 online, It's horrible. I only understood one unit of it. So trying to get the best grade before the final that i will probably fail xD

OpenStudy (anonymous):

how the hell is probability "algebra" forget i asked

OpenStudy (anonymous):

haha there is a lot of things that i don't think should be in an algebra course that were in it

OpenStudy (anonymous):

I totally agree -.-